Mr. x is 62 years old and after 3 years he will be 5 times of his son’s age. What is his son’s age at present?
Answer by mananth(16949)

sons age now = x
fathers age = 62
after 3 years sons age = x+3
fathers age = 65
5(x+3)=65
5x+15=65
5x=65-15
5x=50
/5
x=10 son's age